successive discount of 20% and15% are equal
to a single discount of?

Answer by RichardMathsTutor(3)

discount of 20% = 1-0.2 = 0.8
discount of 15% = 1-0.15 = 0.85
successive discounts of 20% and 15% = 0.8*0.85 = 0.68 = 32% discount